Derivation of the Aggregate Demand Curve Suppose the economy of Y is described by the following...

  1. Derivation of the Aggregate Demand Curve

Suppose the economy of Y is described by the following equations:

Consumption: C = 750 + 0.60 Yd

where Ydrefers to disposable (post-tax) income.

Taxes: T = 300

Government Expenditure: G = 30+0.2Y

Investment: I = 400 -2000r

Money Demand: L(r,Y) = Y – 10,000r Nominal Money Supply :Ms=$12000 Price Level P1=$3

  1. Calculate the tax multiplier in Y
  1. Derive(sketch) the IS curve for Y.
  1. Derive(sketch) the LM curve for Y
  2. Solve for the equilibrium interest rate and output in Y.
Answer #1

Y= C+It9 Y= 750 t0.6 Ya + 400 -2000rt 30+0 +6.21 Y= 750 +0.6(Y-300) +400-2000n +30 + o.2y 0.6Y - (0.6) (300) + 400 - 2000 2 t%3D 12,000 Y- Jo,00 n 3. 4,000 Y - 10,000, 2t wsing Y = 500o 10, o00 n 5000 - 10,000 -10, 600n 4,000= 20,000m TL= 10,000 Y= (4000 5000 10